OpenStudy (rainbow_dash):

A particular sound wave can be graphed using the function . Find the amplitude and period of the function.

OpenStudy (anonymous):

I don't know about you, but I'm willing to bet that this function, whatever this is, looks something like this: \[\huge a\sin \ b\theta\]or\[\huge a\cos \ b\theta\] Whichever it is, your amplitude is going to be given by \[\large |a|\] And your period is given by \[\huge\frac{2\pi}{|b|}\]

OpenStudy (rainbow_dash):

Thanks :)

OpenStudy (anonymous):

Anytime :)